Water can be made to decompose into its elements by using electricity according to the following equation: 2H2O(1) —> 2H2(g) + O2(g)
If you wanted to produce 72g of O2 how many grams of water would you need to decompose? Please explain how/why you got the answer.
1
Expert's answer
2020-12-03T14:01:28-0500

2H2O(1) —> 2H2(g) + O2(g)

n(O2) = m(O2)/Mr(O2) = (72 g)/(32 g/mol) = 2.25 mol

n(H2O) = 2*n(O2) = 2*(2.25 mol) = 4.5 mol from stoichiometric coefficients

m(H2O) = n(H2O)*Mr(H2O) = (4.5 mol)*(18 g/mol) = 81 g.